OpenStudy (appleduardo):

whats the integral for dx/x^3 ?

OpenStudy (anonymous):

The integral? Do you mean the anti-derivative? Remember that \(1/x^3 = x^{-3}\) and \[ \int x^n dx = \frac{x^{n+1}}{n+1}+C,\quad n\ne-1 \]

OpenStudy (appleduardo):

so i'll get: x^-3 and then: \[\frac{ x ^{-2} }{ -2 } \] is that ok?

OpenStudy (anonymous):

That is okay, but so is \[ \frac{x^{-2}}{-2}+1 \]as well. So you're excluding a possible anti-derivative out. That's mean you know.

OpenStudy (anonymous):

Why won't you give \[ \frac{x^{-2}}{-2}-60 \]a chance either? \(-60\) has a family to feed as well.

OpenStudy (anonymous):

@appleduardo ?

OpenStudy (appleduardo):

u mean that i forgot to write aside " + c " :$

OpenStudy (anonymous):

Yeah.

OpenStudy (appleduardo):

:p thank you ! :d